Anle Corporation has a current stock price of $ 18.07 and is expected to pay a dividend of $0.80 in one year. Its expected stock price right after paying that dividend is $20.13.

a. What is​ Anle's equity cost of​ capital?

b. How much of​ Anle's equity cost of capital is expected to be satisfied by dividend yield and how much by capital​ gain?
Answer #1

a.Cost of capital=(End value-Beginning value+Dividends)/Beginning value

=(20.13-18.07+0.8)/18.07

=15.83%(Approx).

b.Dividend yield=Dividend/Beginning value

=(0.8/18.07)=4.43%(Approx).

Capital gains yield=(End value-Beginning value)/Beginning value

=(20.13-18.07)/18.07

=11.40%(Approx).